# Impedance vs. Power Side-channel Vulnerabilities: A Comparative Study

### v1

@@ III-B2 Leakage Model
The HW leakage model considers the HW of the targeted intermediate value as the leakage source. It assumes a linear relationship between the HW and the observed leakage. Let $L$ be the observed leakage, $HW(m)$ be the hamming weight of the targeted intermediate $m$, $a$ and $b$ be the constant coefficients, and $n$ be the noise in the leakage measurement. The observed leakage, $L$, can be expressed as in 4. The fundamental assumption here is that variations in the physical properties observed through side channels during cryptographic operations can reveal information about the internal data being processed.
@@ III-C Threat Model
During the experiment, we write the intermediate values to the memory. It is assumed that the attacker possesses physical access to the system and can control the input plaintexts to the system. The assumptions for the experiment are as follows.
  • In the power side channel case, the attacker can access the PDN of the memory module and measure the power signals when the data is written to the memory using an oscilloscope.
  • In the impedance side channel case, the attacker can connect a VNA to the PDN of the memory module and measure the impedance after each intermediate value corresponding to each plaintext is written to the memory.
@@ III-D Evaluation Metric
To compare the performance of the power and impedance side channels, we use three comprehensive metrics. These metrics assist in the systematic analysis of the effectiveness of both power and impedance side channels in recovering the AES key.
In the evaluation of the two targeted side channels, we use correlation ratio (CR) 5 as a confidence metric for correct key recovery. It is calculated as the ratio between the correlation coefficient for the correct key and the highest correlation coefficient observed for the wrong keys. It can assess the confidence of the correct key guess with respect to other wrong key guesses. A higher ratio indicates a strong level of confidence in precisely guessing the correct value for the key. Let $k_{c}$ and $k_{w}$ denote the correct subkey and the wrong guessed subkey, respectively, from $K$ key space. We use 2 to develop 5 and compute the CR as of the confidence score of each subkey.
@@
In the evaluation for assessing the effectiveness of side-channel attacks, we utilize the interquartile range (IQR) as our third metric. The IQR serves as a robust measure for outlier detection within datasets, which is essential for ensuring the integrity of the data under analysis. The IQR, as defined in 6, calculates the difference between the 75th percentile (Q3) and the 25th percentile (Q1) of the observed data points. This calculation focuses on the variability within the central 50% of the data and renders the IQR a pivotal tool in identifying significant deviations in the correlation coefficients calculated using the hypothetical keys to find the correct key.
The outliers of the correlation coefficient distribution formed by all guessed keys represent the potential correct key candidates. Let $x$ be the calculated correlation coefficient using 2 due to the guessed subkey $K_{x}$. By establishing lower and upper bounds for outlier detection, a correlation coefficient $x$ outside these thresholds is flagged as an outlier, as presented in 7. These outliers represent the potential correct subkey candidates that do not follow the distribution of the wrong subkeys. The incorporation of the IQR metric enhances the evaluation process by facilitating the identification of the correct subkey candidate. This, in turn, allows for a more accurate assessment of side-channel vulnerabilities.
@@ IV-A Hardware Setup
The hardware setup is illustrated in Figure 2. We use an oscilloscope and a vector network analyzer (VNA) to measure the power and impedance signals, respectively. For power measurements, we employ the SIGLENT SDS5104X oscilloscope with a 5 Giga sample per second rate, enabling us to collect 2501 voltage sample points for each measurement. We utilize the Rigol RSA5032N spectrum analyzer for impedance measurements in VNA mode. We select a frequency range from 100 kHz to 3.2 GHz, allowing us to measure impedance at 10,001 linearly distributed frequency points.
Regarding the memory component, we use the FM18W08 FRAM (Ferroelectric Random Access Memory) [34] from Infineon Technologies in this experiment. The FM18W08 is a $32K\times 8$ nonvolatile memory capable of read and write operations like a standard SRAM. It consists of 32,768 memory locations, each with 8 data bits, which can be accessed through a parallel interface. The FRAM array is structured into 4092 rows, each comprising 8 bytes, with addresses of a 15-bit length. We use the Alchitry Au with an Artix 7 FPGA [35] to develop the memory controller to control the FRAM read and write operations. The Alchitry Au operates with a clock frequency of 100 MHz.
@@
Impedance Side-Channel Attack: For the impedance side-channel analysis, we connect the VNA to the 3.3V PDN of the memory chip via a coaxial cable. Unlike in the power side-channel analysis, we collect impedance measurements after the intermediate values have been written to the memory. The VNA measures the reflection coefficient parameters ($S_{11}$), which we convert to impedance values using 1. Although impedance includes both phase and magnitude, we focus solely on the magnitude in our analysis as it carries the most information [9].
@@ IV-B Data Collection and Analysis
For power side-channel measurements, we automate the data collection process by utilizing the write-enable signal of the memory as a trigger. This trigger signal initiates the oscilloscope measurement. We develop a custom Python script incorporating Standard Commands for Programmable Instruments (SCPI) commands to transfer the measured traces to a computer. To mitigate noise, we take each measurement trace as an average of 128 measurements. We use these averaged traces for the power side-channel analysis.
During the impedance side-channel measurements, a similar automated setup is implemented. Python scripts with SCPI commands facilitate the transfer of measured traces from the VNA to the computer. Similar to the power side-channel measurements, we take the average of 100 measurements as each measurement trace to reduce noise.
We perform correlation analysis with the collected measurement traces by comparing the actual leakages with hypothetical leakages. Figure 3 illustrates the analysis steps. We recover the correct key as the key value corresponding to the maximum correlation coefficient observed during the analysis.
V Results and Analysis: Power vs. Impedance
We use the AES-128 encryption and follow the evaluation metrics described in Section III-D to conduct a comprehensive comparison of the performance of power and impedance side channels. Additionally, in a subsequent scenario, we evaluate the resilience of these side channels in a noise-injected environment. In every scenario, we target the first round of the AES encryption process. We adopt a divide-and-conquer approach by dividing the full 128-bit AES key into 16 subkeys and focus our efforts on recovering each subkey individually.
@@ V-A Baseline Side-channel Analysis
To initiate our analysis, we perform the correlation analysis described in Section II-D for all possible 8-bit guessed subkeys and plot the findings in Figure 4. Since we use the 128-bit AES encryption, our analysis involves 16 subkeys. Maximum correlation values in Figure 4 are represented by asterisk (*) for visual clarity. Figure 4(a) represents the results obtained from the power side-channel analysis, while Figure 4(b) depicts the results of the impedance side-channel analysis. The results demonstrate the vulnerability of the 128-bit AES implementation to both the power and impedance side-channel attacks.
Figure 5: Maximum correlation coefficient without noise.
Subkey 0x25 0x0A 0x47 0x38 0xB0 0x52 0x15 0x15
@@
Table I provides a summary of the maximum correlation coefficient observed in both side channels for each individual subkey. Additionally, we compare the maximum correlation coefficient values obtained in Figure 4 for each subkey. Figure 5 presents a visual representation of Table I. The findings show that the maximum correlation values obtained through impedance side channels are consistently higher than those of the power side channels. This suggests a potential superiority of the impedance side channel in information leakage over the power side channel.
Correlation Ratio: We proceed to compare and analyze the results obtained from the two side channels using the metrics discussed in Section III-D. First, to assess the discriminability level of the recovered correct key concerning other key candidates, we calculate the correlation ratio for each subkey using 5. The resulting graph is presented in Figure 6. The findings further suggest that the correct subkeys can be identified with a higher discriminability level most of the time when using the impedance side channel compared to the power side channel.
@@
Our observations indicate that the MTD performs better in the impedance side-channel than the power side-channel. We present the MTD for subkey-1 in Figure 7 for only 128 traces. We find that with this limited number of traces, the correlation coefficient of the correct key is distinctly identified from the wrong key guesses using the impedance side-channel. In contrast, in the case of the power side-channel, the correlation coefficient of the correct key falls within the range of correlation coefficients of the wrong key guesses. We observe similar findings for other subkeys as well. The results suggest that the correct key can be recovered with fewer measurements using the impedance side channel compared to the power side channel.
@@
IQR: To better understand the results and quantify the confidence in the recovered keys, we employ the IQR method on the highest five correlation coefficients obtained from Figure 4. This approach helps us identify the guessed key for which the corresponding correlation coefficient is an outlier. Such an outlier indicates that the correlation coefficient can be statistically distinguished due to that guessed key, making it a potential correct subkey candidate. In scenarios where we find multiple outliers, all of them are considered potential correct subkey candidates. However, the absence of outliers implies that we cannot reliably identify the correct key based on the correlation coefficient values, indicating poor performance.
Subkey 0x25 0x0A 0x47 0x38 0xB0 0x52 0x15 0x15
@@
Table II presents the correlation coefficients that lie outside the first and third quartile range. The entry ‘F’ in the table signifies a failure to distinguish the correct subkey from the wrong ones. The findings suggest that the impedance side channel provides more distinct outliers compared to the power side channel, further corroborating the superiority of the impedance side channel in extracting the correct subkeys.
@@ V-B Noise-Injected Side-channel Analysis
Following the LFSR implementation, we collect the same number of traces as previously for each power and impedance side channel setup and repeat the correlation analysis for AES-128 key extraction. Figure 8 presents the subkey extraction results of each side channel with the LFSR implemented in the background. As illustrated, the performance of the power side channel in extracting the subkeys degrades significantly in the presence of noise. However, even under noisy conditions, the impedance side channel can successfully extract all the 8-bit subkeys of AES-128.
@@
Figure 8: Results of 8-bit subkey extraction for 128-bit AES with noise.
@@ V-C Summary of Results
In synthesizing the findings from this comparative analysis, several key insights emerge regarding the efficacy and implications of impedance and power side-channel attacks in cryptographic systems.
  • Impedance side-channel analysis demonstrates superior efficacy in key extraction, both in the presence and absence of noise, suggesting its potential as a significant security and privacy threat surpassing that of power leakage.
  • Some countermeasures effective against power side-channel attacks may not offer the same level of protection against impedance side-channel attacks. However, we need further example cases and analysis to generalize such a statement.
VI Conclusion
@@
This study provides a comprehensive comparative analysis between impedance and power side-channel analysis techniques, focusing specifically on their efficacy in extracting cryptographic keys. Contrary to the common perception that views impedance side-channel analysis merely as an alternative to power analysis, our findings reveal that impedance analysis holds superior potential in decrypting AES keys where power analysis proves inadequate.
@@
The findings demonstrate the robustness and effectiveness of impedance analysis, challenging its often-overlooked significance. Additionally, the results of this study serve as a compelling motivation for further exploration into the underlying mechanisms of impedance analysis and their implications for hardware security. Furthermore, our findings underscore the need to incorporate impedance side-channel analysis into the standard evaluation protocols for cryptographic systems. By doing so, we can ensure a more comprehensive and robust assessment of potential vulnerabilities, thereby enhancing the overall security posture.

### v3

@@ III-B2 Leakage Model
The HW leakage model considers the HW of the targeted intermediate value as the leakage source. It assumes a linear relationship between the HW and the observed leakage. Let $L$ be the observed leakage, $HW(m)$ be the hamming weight of the targeted intermediate $m$, $a$ and $b$ be the constant coefficients, and $n$ be the noise in the leakage measurement. The observed leakage, $L$, can be expressed as in Eq. 4. The fundamental assumption here is that variations in the physical properties observed through side channels during cryptographic operations can reveal information about the internal data being processed.
@@ III-C Threat Model
During the experiment, we write the intermediate values to the memory. It is assumed that the attacker can control the input plaintexts to the system. The assumptions for the experiment are as follows.
  • In the power SCA, the attacker can access the power delivery network (PDN) of the memory chip and measure the power signals when the data is written to the memory using an oscilloscope.
  • In the impedance SCA, the attacker can connect a VNA to the PDN of the memory module and measure the impedance after each intermediate value corresponding to a plaintext is written to the memory.
@@ III-D Evaluation Metric
To compare the performance of power and impedance side channels, we employ three comprehensive metrics. These metrics do a systematic analysis of the effectiveness of both power and impedance side channels in extracting the AES key.
In the evaluation of the two targeted side channels, we use correlation ratio (CR) Eq. 5 as a confidence metric for correct key recovery. It is calculated as the ratio between the correlation coefficient for the correct key and the highest correlation coefficient observed for the wrong keys. It can assess the confidence of the correct key guess with respect to other wrong key guesses. A higher ratio indicates a strong level of confidence in precisely guessing the correct value for the key. Let $k_{c}$ and $k_{w}$ denote the correct subkey and the wrong guessed subkey, respectively, from $K$ key space. We use Eq. 2 to develop Eq. 5 and compute the CR as of the confidence score of each subkey.
@@
In the evaluation for assessing the effectiveness of side-channel attacks, we use the interquartile range (IQR) as our third metric. The IQR serves as a robust measure for outlier detection within datasets, which is essential for ensuring the integrity of the data under analysis. The IQR, as defined in Eq. 6, calculates the difference between the 75th percentile (Q3) and the 25th percentile (Q1) of the observed data points. This calculation focuses on the variability within the central 50% of the data and renders the IQR a pivotal tool in identifying significant deviations in the correlation coefficients calculated using the hypothetical keys to find the correct key.
@@
The outliers of the correlation coefficient distribution formed by all guessed keys represent the potential correct key candidates. Let $x$ be the calculated correlation coefficient using Eq. 2 due to the guessed subkey $K_{x}$. By establishing lower and upper bounds for outlier detection, a correlation coefficient $x$ outside these thresholds can be flagged as an outlier, as presented in Eq. 7. These outliers represent the potential correct subkey candidates that do not follow the distribution of the wrong subkeys. The incorporation of the IQR metric enhances the evaluation process by facilitating the identification of the correct subkey candidate. This, in turn, allows for a more accurate assessment of side-channel vulnerabilities.
@@ IV-A Hardware Setup
The hardware setup is illustrated in Fig. 2. We use an oscilloscope and a vector network analyzer (VNA) to measure the power and impedance signals, respectively. For power measurements, we employ the SIGLENT SDS5104X oscilloscope with a 5 Giga sample per second rate, enabling us to collect 2501 voltage sample points for each measurement. We utilize the Rigol RSA5032N spectrum analyzer for impedance measurements in VNA mode. We select a frequency range from 100 kHz to 3.2 GHz, allowing us to measure impedance at 10,001 linearly distributed frequency points. Additionally, we use the Alchitry Au development board with an Artix 7 FPGA [37] to develop a memory controller. The memory controller controls the memory read and write operations. The Alchitry Au operates at a clock frequency of 100 MHz.
@@
Impedance Side-Channel Attack: For the impedance SCA, we connect the VNA to the 3.3V PDN of the memory chip via a coaxial cable. Unlike in the power SCA, we collect impedance measurements after the intermediate values have been written to the memory. The VNA measures the reflection coefficient parameters ($S_{11}$), which we convert to impedance values using Eq. 1. Although impedance includes both phase and magnitude, we focus only on the magnitude in our analysis as it carries the most information [9].
@@ IV-B Data Collection and Analysis
For power side-channel measurements, we automate the data collection process by using the write-enable signal of the memory as a trigger. This trigger signal initiates the oscilloscope measurement. We develop a custom Python script incorporating standard commands for programmable instruments (SCPI) commands to transfer the measured traces to a computer. To mitigate noise, we take each measurement trace as an average of 128 measurements. These averaged traces are then used for the power SCA.
During the impedance side-channel measurements, we implement a similar automated setup. Python scripts with SCPI commands facilitate the transfer of measured traces from the VNA to the computer. Similar to the power side-channel measurements, we take the average of 100 measurements as each measurement trace to reduce noise.
We perform correlation analysis with the collected measurement traces by comparing the actual leakages with hypothetical leakages. Fig. 3 illustrates the analysis steps. We recover the correct key as the key value corresponding to the maximum correlation coefficient observed during the analysis.
@@
V Results and Analysis: Power vs. Impedance
We use the AES-128 encryption and follow the evaluation metrics described in Section III-D to conduct a comprehensive comparison of the performance of power and impedance side channels. Additionally, in a subsequent scenario, we evaluate the resilience of these side channels in a noise-injected environment. In every case, we target the S-box substitution process during the first round of AES encryption. We adopt a divide-and-conquer approach by dividing the full 128-bit AES key into 16 subkeys and focus our efforts on recovering each subkey individually.
@@ V-A Baseline Side-channel Analysis
To initiate our analysis, we perform the correlation analysis described in Section II-D for all possible 8-bit guessed subkeys and plot the findings in Fig. 4. Since we use the 128-bit AES encryption, our analysis involves 16 subkeys. Maximum correlation values in Fig. 4 are highlighted by an asterisk (*), and the correlation corresponding to the correct subkey is distinguished by a square ($\square$) for visual clarity. Fig. 4(a) represents the results obtained from the power SCA, while Fig. 4(b) depicts the results of the impedance SCA. The results demonstrate the vulnerability of the 128-bit AES implementation to both the power and impedance side-channel attacks.
Subkey 0x25 0x0A 0x47 0x38 0xB0 0x52 0x15 0xC5
@@
Figure 5: Maximum correlation coefficient without noise.
Table I provides a summary of the maximum correlation coefficient observed in both side channels for each individual subkey. Additionally, we compare the maximum correlation coefficient values obtained in Fig. 4 for each subkey. Fig. 5 presents a visual representation of Table I. The findings show that the maximum correlation values obtained through impedance side channels are consistently higher than those of the power side channels. This suggests a potential superiority of the impedance side channel in information leakage over the power side channel.
Correlation Ratio: We proceed to compare and analyze the results obtained from the two side channels using the metrics discussed in Section III-D. First, to assess the discriminability level of the recovered correct key concerning other key candidates, we calculate the correlation ratio for each subkey using Eq. 5. The resulting graph is presented in Fig. 6. The findings further suggest that the correct subkeys can be identified with a higher discriminability level most of the time when using the impedance side channel compared to the power side channel.
@@
Our observations indicate that the MTD performs better in the impedance side-channel than the power side-channel. We present the MTD for subkey-1 in Fig. 7 for only 128 traces. We find that with this limited number of traces, the correlation coefficient of the correct key is distinctly identified from the wrong key guesses using the impedance side-channel. In contrast, in the case of the power side-channel, the correlation coefficient of the correct key falls within the range of correlation coefficients of the wrong key guesses. We observe similar findings for other subkeys as well. The results suggest that the correct key can be recovered with fewer measurements using the impedance side channel compared to the power side channel.
@@
Subkey 0x25 0x0A 0x47 0x38 0xB0 0x52 0x15 0xC5
@@
IQR: To better understand the results and quantify the confidence in the recovered keys, we employ the IQR method on the highest five correlation coefficients obtained from Fig. 4. This approach helps us identify the guessed key for which the corresponding correlation coefficient is an outlier. Such an outlier indicates that the correlation coefficient can be statistically distinguished due to that guessed key, making it a potential correct subkey candidate. In scenarios where we find multiple outliers, all of them are considered potential correct subkey candidates. However, the absence of outliers implies that we cannot reliably identify the correct key based on the correlation coefficient values, indicating poor performance.
Table II presents the correlation coefficients that lie outside the first and third quartile range. The ‘F’ in the table signifies a failure to distinguish the correct subkey from the wrong ones. The findings suggest that the impedance side channel provides more distinct outliers compared to the power side channel, further corroborating the superiority of the impedance side channel in extracting the correct subkeys.
@@ V-B Noise-Injected Side-channel Analysis
Following the LFSR implementation, we collect the same number of traces as previously for each power and impedance side channel setup and repeat the correlation analysis for AES-128 key extraction. Fig. 8 presents the subkey extraction results of each side channel with the LFSR implemented in the background. As illustrated, the performance of the power side channel in extracting the subkeys degrades significantly in the presence of noise. However, even under noisy conditions, the impedance side channel can successfully extract all the 8-bit subkeys of AES-128.
@@
Figure 8: Results of 8-bit subkey extraction fro 128-bit AES with noise.
@@ V-C Summary of Results
This comparative analysis concludes several key insights regarding the efficacy and implications of impedance and power side-channel attacks within cryptographic systems:
  • Impedance SCA demonstrates superior efficacy in key extraction, demonstrating its potential as a significant security and privacy threat that may surpass that of traditional power leakage attacks, both in noisy and noise-free environments. This capability underscores the importance of addressing impedance-based vulnerabilities in cryptographic systems.
  • Some countermeasures that can mitigate power side-channel attacks (e.g., background noise injection) may not provide equivalent protection against impedance side-channel attacks. However, further case studies and analyses are required to generalize this statement effectively. By exploring a diverse range of scenarios and configurations, we can better understand the specific conditions under which impedance side-channel attacks may succeed or be thwarted. This comprehensive approach will enable us to identify patterns and establish more definitive conclusions regarding the limitations of existing countermeasures against these emerging vulnerabilities.
VI Discussion on the Comparison
In our comparative analysis between impedance-based SCA and power-based SCA, we observed that impedance SCA consistently outperforms power SCA in terms of signal strength and data extraction accuracy. One of the key advantages of impedance SCA is its ability to precisely control the device’s clock, allowing the clock to be paused at specific intervals during impedance measurement. This enables the isolation of switching noise that typically arises from unrelated background activities, ensuring that the observed signal is free from interference. In contrast, power SCA relies on measuring dynamic power consumption, which primarily tracks the flow of current during operation. This method captures not only the switching activity of the desired operation but also the background noise generated by other concurrent switching events within the circuit. The inability to filter out this extraneous noise results in a less distinct signal, making power SCA more susceptible to inaccuracies and noise contamination.
Additionally, impedance measurements are more sensitive to subtle variations in the physical properties of the device, such as material characteristics or component wear, which can provide more nuanced insights into the device’s behavior. Power consumption, by comparison, offers a broader and less differentiated view of the device’s operation, making it harder to isolate specific activities or transitions. Additionally, impedance SCA benefits from lower bandwidth requirements compared to power SCA, as the impedance measurement process does not need to capture the fast transient currents that power SCA requires. This reduced bandwidth requirement simplifies the data acquisition process and enhances the accuracy of the collected data by limiting high-frequency noise.
Importantly, impedance side-channel vulnerabilities remain exploitable even when the chip is powered off, particularly in the context of non-volatile memory. In such cases, the impedance characteristics associated with stored data differ and can be measured without the need to power up the device. This occurs because the impedance of a chip, influenced by specific stored data, does not rely on the device’s power state or power consumption characteristics. This presents a significant challenge for traditional power-based countermeasures, such as power gating. While power gating effectively mitigates power SCA risks by shutting down power to specific sections of a chip during idle periods, it fails to address the vulnerabilities inherent in impedance measurements. As a result, a more comprehensive security approach is necessary.
VII Conclusion
@@
This study provides a comprehensive comparative analysis of impedance and power SCA techniques, focusing specifically on their efficacy in extracting cryptographic keys. Contrary to the common perception that views impedance SCA merely as an alternative to power analysis, our findings find that impedance analysis holds better potential in extracting cryptographic keys, including scenarios where power analysis proves inadequate.
@@
The findings demonstrate the robustness and effectiveness of impedance analysis, challenging its often-overlooked significance. Additionally, the results of this study serve as a compelling motivation for further exploration into the underlying mechanisms of impedance analysis and their implications for hardware security. Furthermore, our findings underscore the need to incorporate impedance SCA into the standard evaluation protocols for cryptographic systems. By doing so, we can ensure a more comprehensive and robust assessment of potential vulnerabilities, thereby enhancing the overall security posture.